OpenStudy (anonymous):

Which is the side length of a cube with a volume of 64 m3?

OpenStudy (anonymous):

\[Cube = m ^{3}\] \[\sqrt[3]{cube} = m\]

Directrix (directrix):

As before, take the cube root. https://www.google.com/?gws_rd=ssl#q=cube+root+of+64

OpenStudy (mathstudent55):

\(64 = 8 \times 8 = (2 \times 2 \times 2) \times (2 \times 2 \times 2) = \) \(= 2 \times 2 \times 2 \times 2 \times 2 \times 2 = (2 \times 2) \times (2 \times 2 )\times (2 \times 2)\) \(= 4 \times 4 \times 4\) \(= 4^3\) \(V = s^3 = 4^3\) \(\color{red}{s}^3 = \color{red}{4}^3\) \(s = 4\)